OpenStudy (mm3591):

Combine 12x+6y+120 and 4x+y=30 to eliminate one of the variables and then solve for the other

OpenStudy (mm3591):

16x+7y=150

OpenStudy (mm3591):

how?

OpenStudy (anonymous):

Hint: Multiply both sides of 4x+y=30 by -3 and then add the equations together.

OpenStudy (anonymous):

12x+6y=120 -12x-3y=-90

OpenStudy (anonymous):

Looks like a typo error. 12x+6y+120 should be 12x+6y=120